OpenStudy (anonymous):

x^2(5x^3-2)^3 dx on [2,0] find the exact value of the definite intergral

OpenStudy (lgbasallote):

hmm this is pretty straightforward \[\huge \int_0^2 x^2 (5x^3 - 2)^3dx\] let u = \((5x^3 - 2)\) du = \(15x^2dx\) \[\huge \implies \frac{1}{15} \int _0^2 u^3 du\] does that help?

OpenStudy (anonymous):

yes a little.

OpenStudy (lgbasallote):

well now use power rule

OpenStudy (lgbasallote):

then plug in the limits and you can be on your merry way

OpenStudy (anonymous):

@hooverst lgbasallote did good work ..just forgot to change limits as well so when u=5x^3-2 when x=0 then u=5(0)^3-2=-2 and when x=2 then u=5(2)^3-2=38 so the integral becomes \[\Large \frac{1}{15}\int\limits_{-2}^{38}{u}^3du\]

OpenStudy (lgbasallote):

hmm...i really dont care about the limits =_= i tend to solve for u..change back to x then just plug in limits..it's less confusing lol

OpenStudy (anonymous):

gud .. but it gets complex little bit the n!! \[\Large \frac{(5x^3-2)^4}{4}\] then limits omg ! :P what about just \[\Large {u^4}\] and then limits

OpenStudy (lgbasallote):

you have your preferences i have mine :p

OpenStudy (anonymous):

ok :P